the sum of the digits of a three digit number is 9. if the digits are reversed the number increases by 495. Find the number if the sum of the tens and hundreds digits is half the units digit.







Answer by lwsshak3(11628)

the sum of the digits of a three digit number is 9. if the digits are reversed the number increases by 495. Find the number if the sum of the tens and hundreds digits is half the units digit.
let u=units digit
let t=tens digit
let h=hundreds digit
..
h+t+u=9
h+t=u/2
(u/2)+u=9
3u/2=9
3u/18
u=6
original number:
100h+10t+6
..
reversed number:
600+10t+h
..
reverse-original number=495
600+10t+h-(100h+10t+6)=495
594-99h=495
99h=99
h=1
t=9-h-u=9-7=2
original number:100h+10t+6=126
reversed number=621
check:621-126=495